OpenStudy (anonymous):

What is the magnitude and direction of if G(2, -2) and H(7, 6)? Answer magnitude: 9.43 units; direction 32.01° magnitude: 9.43 units: direction: 57.99° magnitude: 6.4 units: direction: 32.01° magnitude: 6.4 units; direction: 57.99°

OpenStudy (anonymous):

my bro is here to help me

OpenStudy (anonymous):

Distance formula gives you the magnitude

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So \[\sqrt{(7-2)^2 + (6-(-2))^2}\]

OpenStudy (anonymous):

\[\sqrt{5^2 + 8^2}\]

OpenStudy (anonymous):

& for direction use tan(angle )

OpenStudy (anonymous):

25 + 64

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so \[\sqrt{89}\]

OpenStudy (anonymous):

Im not sure what that number is. I cant find my calculator.

OpenStudy (anonymous):

Its 9.something

OpenStudy (anonymous):

its 9.43

OpenStudy (anonymous):

Ok good.

OpenStudy (anonymous):

what about the second half

OpenStudy (anonymous):

Ok lemme see, i forgot how to do that.

OpenStudy (anonymous):

i dot to go sorry thank u ill get it later

OpenStudy (anonymous):

Ok sorry i forgot how to do the second part.

OpenStudy (anonymous):



OpenStudy (anonymous):

You need to measure this angle: